62. (x – 10)(x – 2) = -20 thanks for all your help, i really am trying
Answer by ptaylor(2198)

Lets first get the equation into quadratic form (ax^2+bx+c=0) and we can then apply the quadratic formula x =(-b +- sqrt( b^2-4ac ))/2a to get the two solutions.
Expanding the left side of the equation (eg., multiplying (x-10)(x-2)), we get x^2-12x+20. We now have:
x^2-12x+20=-20, adding 20 to both sides, we have:
x^2-12x+40=0 It's now in quadratic form and we can apply the quadratic formula to get the answer
x=(12+-sqrt(144-160))/2 or
x=(12+-sqrt(-16))/2 ; which yields imaginary solutions:
x=6+2i and x=6-2i
Hope this helps-----ptaylor